# Monocular Lane Detection Based on Deep Learning: A Survey

### v1

@@ III-B2 Grids-based modeling
It is obvious that $h\ll H$ and $w\ll W$. In this way, the original pixel by pixel classification requires a time complexity of $O(H\times W\times C)$, while the time complexity of grid classification is $O(h\times w\times C)$, where C is the number of classes. Then, a large fully connected (FC) layer is used to output the classification probability for each grid. This approach leverages global features as input, eliminating the need for upsampling as required in semantic segmentation. As a result, it achieves a larger receptive field, which is beneficial for handling lane occlusion and varying lighting conditions. Additionally, a lane structure loss is designed based on the thin and elongated prior characteristics of lanes. The proposed network not only achieves strong performance but also enables ultra-fast inference.
@@
Figure 5: More detial of different lane modeling. (a) A lane in FV image. (b) Mask-based modeling, which perform semantic segmentation. (c) Grids-based modeling, which execute the row/column-wise classification. (d) Keypoints-based modeling, which detect discrete points on the lane. (e) Line Anchor-based modeling, which learn the horizontal offset of equidistant points. (f) Curve-based modeling, which predict the curve parameters.
The strategy of row-wise classification leverages the vertical, slender nature of lanes. Unfortunately, it is not well-suited for some curved or near-horizontal lanes, where several meshes in a row may correspond to the same lane. For this reason, UFLD-V2 [75] extends row-wise classification to row/column-wise classification to address the issue that row-wise classification cannot handle horizontal lanes. However, it still employs a multi-classification strategy [36] for lane instances, which result in an overly simplistic choice between row and column classification, thereby limiting its generalizability in real-world scenarios. CANet [76] further optimizes this approach. It employ U-shaped guidelines to constrain lane instance kernel generation on the basis of CondLaneNet [17]. An adaptive decoder is designed, which dynamically chooses between row-by-row or column-by-column classification for each instance.
Since the networks output the classification probability of lanes in each row/column grid, the post-processing is also required to obtain the final vectorized result. Specifically, each point’s coordinate is calculated as the expectation of locations (grids from the same row/column), i.e. a weighted average by probability.
TABLE II: Summary of representative 2D lane detection methods. Under Task Paradigm, ”Seg” for segmentation-based methods, ”ODet” for object detection-based methods. For segmentation-based methods, the manner of instance discrimination is additionally indicated instead of using ”✓” to mark. $\uparrow$” and $\downarrow$” represent the instance discrimination via bottom-up and top-down approach in segmentation-based methods, respectively. ”Max. L.” represents predefining maximum number of lanes, and ”Dy. K.” means that predicting dynamic kernels.
@@
Seg ODet (Solution for Obscured Lanes)
@@ III-B3 Keypoints-based modeling
The mask-based modeling methods often involves predicting numerous irrelevant regions, so some research efforts attempt to directly predict keypoints of lanes. This, like grids-based modeling, can be seen as a sparse version of mask-based modeling, but it directly provides the vectorized expression required downstream.
Some works follow a bottom-up approach. PINet [78] uses a stacked hourglass network to predict keypoint locations and feature embeddings, and clusters different lane instances based on the similarity of feature embeddings. FOLOLane [67] estimates existence and offset of local lane keypoints, and designs a decoder module with low-level operators integrates the local information into curve instances. Only keypoints on adjacent boundaries are paired, allowing the network to better focus on detailed features, but the lack of global features causes poor performance under the obscured lanes scene. To add global information, GANet [18] adopts a more efficient post-processing method to cluster points by directly calculating the offset between the keypoint and the start point to globally return to the keypoint. Additionally, a lane-aware feature aggregator based on deformable convolution (DCN) [90] is proposed to improve the shape of lanes and better capture the local context on the lane. RCLane [84] sparsifies the binary segmentation results to obtain the keypoints of all lanes. It decodes the channel in a chained mode using distance and transmission head to predict the keypoints and their continuous relationships in the channel. It also proposes a bilateral prediction method for learning complex topology and global shape information, which can adapt to lanes with complex structures, such as Y-shaped and forked lanes. LanePtrNet [91] designs a centrality farthest point sampling method to determine the lane center point. Then a grouping head performs clustering based on the center point position and lane point embeddings to obtain the final lane.
Others adopt a top-down manner. For each lane, Chougule et al. [92] directly regress the position of keypoints, and Yoo et al. [93] predict a feature map and searches for keypoints of the lane on each row. However, both of them distinguish instances using a multi-classification strategy, as described in Section III-B1, which is not flexible. CondLSTR [69] improves the instance-obtaining method in CondLaneNet [17]. It leverage Transformer to generate dynamic and offset kernels for each lane, enhancing global knowledge. Then, these kernels are dynamically convolved with the entire feature map to predict the heatmap and offset maps of lane keypoints.
@@ III-C1 Line Anchor-based modeling
A lane on the image can be represented by equally-spaced 2D-points. Specifically, lane is expressed as a sequence of points, i.e., $P={(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})}$. The y-coordinate of points is eaually sampled through image vertically, i.e., $y_{i}=\frac{H}{N-1}\cdot i$, where $H$ is the height of image. Accordingly, the x-coordinate is associated with the respective $y_{i}\in Y$. With $P$ and $y_{i}$, the positions of points that form a lane can be located.
@@
Line-CNN [77] uses a large number of predefined straight lines as line anchors. However, it predicts the scores, lengths, and transverse coordinate offsets of all anchors based on the local features of each start point, which implies that the feature map from the backbone must have a sufficiently high number of channels. To slove this problem, LaneATT [80] proposes a line anchor feature pooling method that allows the use of a lightweight backbone and presents a high-performance and efficient attention aggregation mechanism to better detect Obscured lanes. PointLaneNet [94] and CurveLane-NAS [40] separate images into non-overlapping grids and regress lanes based on vertical line anchors. Especially, CurveLane-NAS uses Network Architecture Search (NAS) [95] to find a better network to capture more accurate information, which is beneficial for detecting curved lanes. SGNet [81] introduces a novel vanishing-point oriented anchor generator and adds multiple structural guides to improve performance. Jin et al. [43] introduce data-driven descriptors called eigenlanes. They obtain line anchors that can better regress curved lanes using lower-order approximations of the lane matrix. Non-maximum suppression (NMS) post-processing is also unavoidable due to the limitations of a large number of predefined anchors and early positive and negative sample matching strategies in object detection.
@@
With the widespread application of Transformers in object detection, research in this field gradually shifts from dense prediction paradigms, such as YOLO [96] and Faster R-CNN [97], to set prediction paradigms like DETR [98]. Similarly, lane detection based on object detection shifts from a fixed dense approach to a dynamic sparse approach. Based on Deformable DETR [99], Laneformer [83] introduces two novel row and column self-attention operations in the encoder to effectively capture lane context. The binary matching strategy enables an NMS-free approach. Inspired by Sparse R-CNN [100], CLRNet [19] uses multi-scale feature maps at the pyramid level to iteratively adjust the positions of a small set of preset line anchors [101]. It presents ROIGather to enhance lane feature extraction, effectively addressing challenges such as occlusion and lighting variations. Additionally, Line IoU Loss is introduced for global lane regression, enhancing positioning accuracy. CLRNet achieves state-of-the-art results on 2D lane detection datasets. The improvements to Line IoU Loss and label matching are made in [102] and [103], respectively. O2SFormer [104] proposes a one-to-many label allocation strategy and incorporates lane anchor points into position queries [105], providing explicit positional priors that accelerate model convergence. ADNet [86] removes the limitation of anchor starting points by learning heatmaps for these points and their related directions, enabling the network to adapt to diverse lane types across different datasets. It puts forward a module based on a hybrid CNN&Transformer architecture [106] [107] to expand the receptive field, and proposes the Generalized Line IoU Loss to address the limitations of Line IoU Loss [19]. Similarly, SRLane [87] generates sparse line anchors by predicting local directional heatmaps and develops a lane segment association module to adjust non-fitting lane segments. Sparse Laneformer [108] designs learnable lane and angle queries to generate sparse line anchors. It employs a two-stage Transformer decoder to refine lane predictions. GSENet [89] designs a global feature extraction module based on dilated convolution [51] and SimAm [109] to obtain accurate and comprehensive global features, and further enhances semantic representation using ViT [110]. HGLNet [88] leverages the large receptive field of dilated convolution to enhance the representation of local features, and designed a global extraction head based on deformable-attention [99] to extract global feature of lanes adaptively.
@@ III-C2 Curve-based modeling
Several studies model lanes as curve equations in image space, predicting the parameters of the modeled curves. A pioneering work [111] proposes a differentiable least squares fitting module, which fits the polynomial curves (e.g. $x=ay^{3}+by^{2}+cy+d$) to the points predicted by deep neural networks. This approach mainly consists of two components: a deep network for predicting weight maps and a differentiable least-squares fitting module to generate the best fit curve. Then, PolyLaneNet [79] directly learns to predict polynomial coefficients with simple fully connected layers. LSTR [82] use transformer to predict polynomials in an end-to-end manner based on DETR [98]. However, the performance of these methods remains suboptimal due to the difficulty of curve’s parameter learning and challenges in transformer training. PGA-Net [85] introduces an improved supervised strategy to accelerate transformer convergence and proposes a Mean Curvature Loss to constrain the curvature of predicted lanes, enhancing the predictive accuracy for curved lanes.
Feng et al. [20] argue that polynomial are abstract and the coefficients are challenging to optimize, recommending third-order Bézier curves for lane modeling. The method predicts four control points to determine lane positions, proving more robust than direct regression of polynomial coefficients. They consider the pseudo-symmetry of lanes in images and propose a feature-flipping fusion module based on DCN [112] to enhance feature representation in vehicle front-view images. Subsequently, Chen et al. [113] model lanes as more flexible B-spline curves and propose a novel curve-distance calculation method to improve control points prediction supervision.
@@
Finally, substituting Eqn. 7 into Eqn. 6 can obtain the ego vehicle coordinates. Clearly, it is cumbersome to transform every pixel coordinate according to this manner. We can select four points on FV as Region Of Interest (ROI) and calculate the corresponding positions in the ego vehicle coordinate system using the above method. Then we can establish a system of ternary linear equations:
@@ IV-A Classification Framework
As shown in Figure 7, existing 3D lane detection methods can be divided into two categories: (a) BEV-based method, which utilize camera parameters and convert the extracted FV features into BEV features with height information in some way. This process of constructing an intermediate proxy is usually referred to as view transformation [117]. Through this way, the 3D lane detection task can be simplified to 2D lane detection in BEV, and then combining it with the corresponding height values estimated by a height estimation head yields the final three-dimensional lanes. Therefore, the performance of this type of method depends not only on the 2D lane detection results in BEV but also on the adopted view transformation method. (b) BEV-free methods, which do not hinge on BEV features. It can be further divided into two types. One is to detect 2D lanes in the FV image while predicting their depth, and then projecting them onto the 3D space. The other is to directly model lanes in the 3D space. With the initialized 3D information, it is possible to project it onto FV based on camera parameters. This approach enables directly interactions between the 3D lane and FV features, ultimately refining and updating the 3D lane.
@@ IV-B BEV-based methods
The pipeline follows the established process of dense BEV perception methods [126, 127, 128]. Since BEV features inherently conceal height information, the subsequent lane decoding process only needs to consider the 2D BEV plane, which can naturally be integrated with 2D lane detection methods. The view transformation between FV and BEV feature can be formulated as:
@@ IV-B1 IPM for View Transformation
3D-LaneNet [22] is the first method which uses deep learning predicting 3D lanes directly from monocular images. The network first predicts the camera pitch angle and height to generate a differentiable IPM, combining the original FV feature map to create the BEV feature map. In the BEV plane, lanes are represented as vertical line anchors. 3D-LaneNet demonstrates promising results in detecting 3D lanes from monocular images. Gen-LaneNet [23] directly uses 2D lane segmentation results as input for IPM, allowing for the utilization of extensive 2D lane data and enhancing the model’s generalization. In contrast to 3D-LaneNet, Gen-LaneNet offers more reliable supervision by using the camera pitch angle and height as GT. Li et al. [118] propose a new loss function based on Gen LaneNet to better extract the height information of 3D lanes from 2D lane representations. In the BEV plane, the above work uses vertical line anchors to model lanes. However, matching line anchors to ground truth is performed by measuring the distance at a predefined $y=Y_{ref}$, which may result in missed detections for short lanes. Therefore, 3D-LaneNet+ [129] avoids this issue using a bottom-up segmentation approach. Liu et al. [130] believe that the model can be independent of the ground truth camera pose provided by the benchmark. They designed a two-stage network based on Transformer, which first predicting camera pose, i.e., the required parameters for IPM, then extracts BEV features, and finally regresses polynomial coefficients. 3D-SplineNet [122] treats lanes as B-spline curves in the BEV plane.

### v2

@@ III-B2 Grids-based modeling
It is obvious that $h\ll H$ and $w\ll W$. In this way, the original pixel-by-pixel classification requires a time complexity of $O(H\times W\times C)$, while the time complexity of grid classification is $O(h\times w\times C)$, where C is the number of classes. Then, a large fully connected (FC) layer is used to output the classification probability for each grid. This approach leverages global features as input, eliminating the need for upsampling as required in semantic segmentation. As a result, it achieves a larger receptive field, which is beneficial for handling lane occlusion and varying lighting conditions. Additionally, a lane structure loss is designed based on the thin and elongated prior characteristics of lanes. The proposed network not only achieves strong performance but also enables ultra-fast inference.
@@
Figure 5: More detail of different lane modeling. (a) A lane in FV image. (b) Mask-based modeling, which performs semantic segmentation. (c) Grids-based modeling, which executes the row/column-wise classification. (d) Keypoints-based modeling, which detects discrete points on the lane. (e) Line Anchor-based modeling, which learns the horizontal offset of equidistant points. (f) Curve-based modeling, which predicts the curve parameters.
The strategy of row-wise classification leverages the vertical, slender nature of lanes. Unfortunately, it is not well-suited for some curved or near-horizontal lanes, where several meshes in a row may correspond to the same lane. For this reason, UFLD-V2 [75] extends row-wise classification to row/column-wise classification to address the issue that row-wise classification cannot handle horizontal lanes. However, it still employs a multi-classification strategy [36] for lane instances, which results in an overly simplistic choice between row and column classification, thereby limiting its generalizability in real-world scenarios. CANet [76] further optimizes this approach. It employs U-shaped guidelines to constrain lane instance kernel generation based on CondLaneNet [17]. An adaptive decoder is designed, which dynamically chooses between row-by-row or column-by-column classification for each instance.
Since the networks output the classification probability of lanes in each row/column grid, post-processing is also required to obtain the final vectorized result. Specifically, each point’s coordinate is calculated as the expectation of locations (grids from the same row/column), i.e. a weighted average by probability.
TABLE II: Summary of representative 2D lane detection methods. Under Task Paradigm, ”Seg” for segmentation-based methods, ”ODet” for object detection-based methods. For segmentation-based methods, the manner of instance discrimination is additionally indicated instead of using ”✓” to mark. $\uparrow$” and $\downarrow$” represent the instance discrimination via bottom-up and top-down approach in segmentation-based methods, respectively. ”Max. L.” represents predefining the maximum number of lanes, and ”Dy. K.” means that predicting dynamic kernels.
@@
Seg ODet (Solution for Obscure Lanes)
@@ III-B3 Keypoints-based modeling
The mask-based modeling methods often involve predicting numerous irrelevant regions, so some research efforts attempt to directly predict keypoints of lanes. This, like grids-based modeling, can be seen as a sparse version of mask-based modeling, but it directly provides the vectorized expression required downstream.
Some works follow a bottom-up approach. PINet [78] uses a stacked hourglass network to predict keypoint locations and feature embeddings, and clusters different lane instances based on the similarity of feature embeddings. FOLOLane [67] estimates the existence and offset of local lane keypoints, and designs a decoder module with low-level operators that integrates the local information into curve instances. Only keypoints on adjacent boundaries are paired, allowing the network to better focus on detailed features, but the lack of global features causes poor performance under the obscure lanes scene. To add global information, GANet [18] adopts a more efficient post-processing method to cluster points by directly calculating the offset between the keypoint and the start point to globally return to the keypoint. Additionally, a lane-aware feature aggregator based on deformable convolution (DCN) [90] is proposed to improve the shape of lanes and better capture the local context on the lane. RCLane [84] sparsifies the binary segmentation results to obtain the keypoints of all lanes. It decodes the channel in a chained mode using distance and transmission head to predict the keypoints and their continuous relationships in the channel. It also proposes a bilateral prediction method for learning complex topology and global shape information, which can adapt to lanes with complex structures, such as Y-shaped and forked lanes. LanePtrNet [91] designs a centrality farthest point sampling method to determine the lane center point. Then a grouping head performs clustering based on the center point position and lane point embeddings to obtain the final lane.
Others adopt a top-down manner. For each lane, Chougule et al. [92] directly regress the position of keypoints, and Yoo et al. [93] predict a feature map and searches for keypoints of the lane on each row. However, both of them distinguish instances using a multi-classification strategy, as described in Section III-B1, which is not flexible. CondLSTR [69] improves the instance-obtaining method in CondLaneNet [17]. It leverages Transformer to generate dynamic and offset kernels for each lane, enhancing global knowledge. Then, these kernels are dynamically convolved with the entire feature map to predict the heatmap and offset maps of lane keypoints.
@@ III-C1 Line Anchor-based modeling
A lane on the image can be represented by equally spaced 2D points. Specifically, the lane is expressed as a sequence of points, i.e., $P={(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})}$. The y-coordinate of points is equally sampled through the image vertically, i.e., $y_{i}=\frac{H}{N-1}\cdot i$, where $H$ is the height of the image. Accordingly, the x-coordinate is associated with the respective $y_{i}\in Y$. With $P$ and $y_{i}$, the positions of points that form a lane can be located.
@@
Line-CNN [77] uses a large number of predefined straight lines as line anchors. However, it predicts the scores, lengths, and transverse coordinate offsets of all anchors based on the local features of each start point, which implies that the feature map from the backbone must have a sufficiently high number of channels. To solve this problem, LaneATT [80] proposes a line anchor feature pooling method that allows the use of a lightweight backbone and presents a high-performance and efficient attention aggregation mechanism to better detect obscure lanes. PointLaneNet [94] and CurveLane-NAS [40] separate images into non-overlapping grids and regress lanes based on vertical line anchors. In particular, CurveLane-NAS uses Network Architecture Search (NAS) [95] to find a better network to capture more accurate information, which is beneficial for detecting curved lanes. SGNet [81] introduces a novel vanish point-oriented anchor generator and adds multiple structural guides to improve performance. Jin et al. [43] introduce data-driven descriptors called eigenlanes. They obtain line anchors that can better regress curved lanes using lower-order approximations of the lane matrix. Non-maximum suppression (NMS) post-processing is also unavoidable due to the limitations of a large number of predefined anchors and early positive and negative sample matching strategies in object detection.
@@
With the widespread application of Transformers in object detection, research in this field gradually shifts from dense prediction paradigms, such as YOLO [96] and Faster R-CNN [97], to set prediction paradigms like DETR [98]. Similarly, lane detection based on object detection shifts from a fixed dense approach to a dynamic sparse approach. Based on Deformable DETR [99], Laneformer [83] introduces two novel row and column self-attention operations in the encoder to effectively capture lane context. The binary matching strategy enables an NMS-free approach. Inspired by Sparse R-CNN [100], CLRNet [19] uses multi-scale feature maps at the pyramid level to iteratively adjust the positions of a small set of preset line anchors [101]. It presents ROIGather to enhance lane feature extraction, effectively addressing challenges such as occlusion and lighting variations. Additionally, Line IoU Loss is introduced for global lane regression, enhancing positioning accuracy. CLRNet achieves state-of-the-art results on 2D lane detection datasets. The improvements to Line IoU Loss and label matching are made in [102] and [103], respectively. O2SFormer [104] proposes a one-to-many label allocation strategy and incorporates lane anchor points into position queries [105], providing explicit positional priors that accelerate model convergence. ADNet [86] removes the limitation of anchor starting points by learning heatmaps for these points and their related directions, enabling the network to adapt to diverse lane types across different datasets. It puts forward a module based on a hybrid CNN&Transformer architecture [106] [107] to expand the receptive field, and proposes the Generalized Line IoU Loss to address the limitations of Line IoU Loss [19]. Similarly, SRLane [87] generates sparse line anchors by predicting local directional heatmaps and develops a lane segment association module to adjust non-fitting lane segments. Sparse Laneformer [108] designs learnable lane and angle queries to generate sparse line anchors. It employs a two-stage Transformer decoder to refine lane predictions. GSENet [89] designs a global feature extraction module based on dilated convolution [51] and SimAm [109] to obtain accurate and comprehensive global features, and further enhances semantic representation using ViT [110]. HGLNet [88] leverages the large receptive field of dilated convolution to enhance the representation of local features. It designs a global extraction head based on deformable-attention [99] to extract global feature of lanes adaptively.
@@ III-C2 Curve-based modeling
Several studies model lanes as curve equations in image space, predicting the parameters of the modeled curves. A pioneering work [111] proposes a differentiable least squares fitting module, which fits the polynomial curves (e.g. $x=ay^{3}+by^{2}+cy+d$) to the points predicted by deep neural networks. This approach mainly consists of two components: a deep network for predicting weight maps and a differentiable least-squares fitting module to generate the best-fit curve. Then, PolyLaneNet [79] directly learns to predict polynomial coefficients with simple fully connected layers. LSTR [82] uses Transformer to predict polynomials in an end-to-end manner based on DETR [98]. However, the performance of these methods remains suboptimal due to the difficulty of the curve’s parameter learning and challenges in transformer training. PGA-Net [85] introduces an improved supervised strategy to accelerate transformer convergence and proposes a Mean Curvature Loss to constrain the curvature of predicted lanes, enhancing the predictive accuracy for curved lanes.
Feng et al. [20] argue that polynomials are abstract and the coefficients are challenging to optimize, recommending third-order Bézier curves for lane modeling. The method predicts four control points to determine lane positions, proving more robust than direct regression of polynomial coefficients. They consider the pseudo-symmetry of lanes in images and propose a feature-flipping fusion module based on DCN [112] to enhance feature representation in vehicle front-view images. Subsequently, Chen et al. [113] model lanes as more flexible B-spline curves and propose a novel curve-distance calculation method to improve control points prediction supervision.
@@
Finally, substituting Eqn. 7 into Eqn. 6 can obtain the ego vehicle coordinates. It is cumbersome to transform every pixel coordinate according to this manner. We can select four points on FV as Region Of Interest (ROI) and calculate the corresponding positions in the ego vehicle coordinate system using the above method. Then we can establish a system of ternary linear equations:
@@ IV-A Classification Framework
As shown in Figure 7, existing 3D lane detection methods can be divided into two categories: (a) BEV-based method, which utilize camera parameters and convert the extracted FV features into BEV features with height information in some way. This process of constructing an intermediate proxy is usually referred to as view transformation [117]. In this way, the 3D lane detection task can be simplified to 2D lane detection in BEV, and then combining it with the corresponding height values estimated by a height estimation head yields the final three-dimensional lanes. Therefore, the performance of this type of method depends not only on the 2D lane detection results in BEV but also on the adopted view transformation method. (b) BEV-free methods, which do not hinge on BEV features. It can be further divided into two types. One is to detect 2D lanes in the FV image while predicting their depth, and then projecting them onto the 3D space. The other is to directly model lanes in the 3D space. With the initialized 3D information, it is possible to project it onto FV based on camera parameters. This approach enables direct interactions between the 3D lane and FV features, ultimately refining and updating the 3D lane.
@@ IV-B BEV-based methods
The pipeline follows the established process of dense BEV perception methods [126, 127, 128]. Since BEV features inherently conceal height information, the subsequent lane decoding process only needs to consider the 2D BEV plane, which can naturally be integrated with 2D lane detection methods. The view transformation between FV and BEV features can be formulated as:
@@ IV-B1 IPM for View Transformation
3D-LaneNet [22] is the first method which uses deep learning to predict 3D lanes directly from monocular images. The network first predicts the camera pitch angle and height to generate a differentiable IPM, combining the original FV feature map to create the BEV feature map. In the BEV plane, lanes are represented as vertical line anchors. 3D-LaneNet demonstrates promising results in detecting 3D lanes from monocular images. Gen-LaneNet [23] directly uses 2D lane segmentation results as input for IPM, allowing for the utilization of extensive 2D lane data and enhancing the model’s generalization. In contrast to 3D-LaneNet, Gen-LaneNet offers more reliable supervision by using the camera pitch angle and height as GT. Li et al. [118] propose a new loss function based on Gen LaneNet to better extract the height information of 3D lanes from 2D lane representations. In the BEV plane, the above work uses vertical line anchors to model lanes. However, matching line anchors to ground truth is performed by measuring the distance at a predefined $y=Y_{ref}$, which may result in missed detections for short lanes. Therefore, 3D-LaneNet+ [129] avoids this issue using a bottom-up segmentation approach. Liu et al. [130] believe that the model can be independent of the ground truth camera pose provided by the benchmark. They designed a two-stage network based on Transformer, which first predicts camera pose, i.e., the required parameters for IPM, then extracts BEV features, and finally regresses polynomial coefficients. 3D-SplineNet [122] treats lanes as B-spline curves in the BEV plane.